# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v1

@@ Abstract.
As topological applications of main results(see Corollory 1.7) we show that for a complete noncompact Riemannian manifold on which the Sobolev inequality (1.8) holds true, $\dim(M)=n\geq 3$ and ${\rm Ric}(x)\geq 0$ outside some geodesic ball $B(o,R_{0})$, there exists a positive constant $C(n)$ depending only on $n$ such that, if
@@ 1. Introduction
The renowned Cheng-Yau’s logarithm gradient estimate showed that when $p=2$, any solution bounded from above or below to (1.4) is a constant provided the Ricci curvature of the Riemannian manifold is nonnegative (cf. [16]). Kotschwar-Ni [35] proved that any positive $p$-harmonic function on a complete Riemannian manifold with nonnegative sectional curvature is a positive constant. Subsequently, this result was verified to remain correct by Wang and Zhang [57] under the condition of nonnegative Ricci curvature, where the authors applied the Nash-Moser iteration technique and Saloff-Coste’s Sobolev inequality (cf. [48, Theorem 3.1]) to study the gradient estimates of equation (1.4). This gradient estimate was later refined to be sharp by Sung and Wang [54].
@@
Especially, in the case $p=2$, $a>0$ and $\alpha\in(-\infty,\,\frac{n+2}{n-2})$, Lu [40] established the Cheng-Yau type logarithm gradient estimate for positive solutions to Lane-Emden equation (1.2) on a complete Riemannian manifold with Ricci curvature bounded from below.
@@
We notice in this article that if the $\chi$-type Sobolev inequality holds on $(M,g)$ with $\mathrm{dim}(M)\geq 3$, then $\chi\leq n/(n-2)$ (see Theorem 2.1). Hence, we assume $\chi\in(1,n/(n-2)]$ when $\mathrm{dim}(M)\geq 3$ in the rest of this article. Clearly, (1.7) is just the well-known Sobolev inequality if $\chi={n}/(n-2)$, i.e.
@@
which holds true for a large class of complete Riemannian manifolds. Indeed, significant efforts have recently been devoted to studying optimal Sobolev inequalities on Riemannian manifolds (see the survey [27] and its references). In particular, Brendle [7] obtained the sharp Sobolev inequality on complete noncompact Riemannian manifolds with nonnegative Ricci curvature using the so-called ABP method, addressing an open question posed by Cordero-Erausquin, Nazaret, and Villani [21] for such manifold. Balogh and Kristály [3] provide an alternative proof of Brendle’s rigidity result and confirmed that the Sobolev constant is sharp. Concretely, their main result can be stated as follows:
$\omega_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $\mathcal{S}\left(\mathbb{R}^{n}\right)$ is the best Sobolev constant in $\mathbb{R}^{n}$.
@@
It was first observed by the first named author [58, Proposition 2.4], and later by Carron (cf. [12]) and Akutagawa (cf. [1]) independently that if the Sobolev constant of a Riemannian manifold $(M^{n},g)$ is positive, then the volume of the geodesic ball of radius $R$ is larger than $CR^{n}$, where the constant $C$ depends only on $n$ and the Sobolev constant.
@@
    In a certain sence, these two theorems can be seen as a effective version of the classical Liouville results due to Gidas-Spruck [29] and Serrin-Zou [51]. To see this, let $g_{\epsilon}$ be a sequence of metrics on $\mathbb{R}^{n}$ with $g_{\epsilon}$ equals to the Euclidean metric outside some compact set of $\mathbb{R}^{n}$ and with $g_{\epsilon}$ converges to the Euclidean metric in $C^{2}$ sense as $\epsilon\rightarrow 0$. Then the Positive Mass Theorem tells us that $g_{\epsilon}$ can not be of nonnegative Ricci curvature. Since the Sobolev inequality only depends on the $C^{0}$ property of the metric. Our results indicate that the classical Liouville property also holds if the deformed metric $g_{\epsilon}$ satisfied
@@
Now, we turn to considering the topological properties of a noncompact complete manifold with nonnegative Ricci curvature outside a compact set. In 1991 Cai [9] showed that such a manifold is of finitely many ends. Later, in 1995 Cai, Colding and Yang [10] discussed the gap phenomenon for ends of such a class of manifold, concretely, they proved the following theorem:
@@
On the other hand, using harmonic function theory to study the number of ends has a long history (see for example, [11, 26, 36]). Especially, if the Sobolev constant of $(M,g)$ is positive and $\dim(M)\geq 3$, the first named author has ever used harmonic function theory to prove that such a manifold is of finitely many ends and the dimension of linear space spanned by bounded harmonic functions on $(M,g)$ equals the number of ends (cf. [58]). As a geometric application of Corollory 1.4, we obtain the following gap theorem for ends.
@@
The rest of our paper is organized as follows. In section 2, we will give the volume estimate of a geodesic ball $B_{R}$ in $(M,g)$ on which a $\chi$-type Sobolev inequality holds. Section 3 is devoted to a meticulous estimate of $\mathcal{L}\left(|\nabla\log v|^{2\alpha}\right)$ (the explicit definition of the operator $\mathcal{L}$is given in (3.3)). The proofs of our main results reveal that, by selecting an appropriate parameter $\alpha$, we can establish effective integral estimates for the gradient of positive solutions to equation (1.1). In particular, when $\rm Ric$$_{-}$ satisfies the condition stated in Theorem 1.2 or Theorem 1.5, we obtain a $L^{\theta\chi}$ bound of $|\nabla\log v|$. The crucial thing is this bound only depends on the volume and the radius of the geometric ball. With this integral bound in hand, we then apply the Nash-Moser iteration scheme to complete the proof of Theorem 1.5. In section 4, we continue to prove Theorem 1.3 by choosing suitable auxiliary functions. In Section 5 we provide the proof of Theorem 1.6.
@@ 2. Volume estimate
We conclude that
@@
$\lim_{m\rightarrow\infty}\alpha(m)=\frac{\chi}{\chi-1}\quad\text{and}\quad\lim_{m\rightarrow\infty}\beta(m)=\frac{\chi(2\chi-1)}{(\chi-1)^{2}}.$
@@ 5. Geometric applications
Using harmonic function theory to study geometric and topological properties of manifolds has a long history. Here we take some examples. Denote the linear space spanned by bounded harmonic functions on $M$ by $H^{\infty}(M)$. The first named author showed the following
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold with Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and ${\rm Ric}(M)\geq 0$ outside some compact subset. Then $M$ has only finitely many ends $E_{1},E_{2},\cdots,E_{k}$ and $\dim H^{\infty}(M)=k$.
@@
For the sake of completeness, here we give the routine to construct bounded positive harmonic functions on such a manifold $(M,g)$ which has at least two ends and the Sobolev constant of $(M,g)$ is of a positive lower bound. Denote the two ends of $(M,g)$ as $E_{\alpha}$ and $E_{\beta}$, and let $\Omega_{i}$ ($i=1,2,\cdots$) be an exhaustion of $(M,g)$, i.e., each $\Omega_{i}$ is an open domain contained in $M$ and $\bar{\Omega}_{i}$ is compact, $\bar{\Omega}_{i}\subset\Omega_{i+1}$ for every $i\geq 1$, and $\cup_{i=1}^{\infty}\Omega_{i}=M$.
@@
Moreover, denote $E_{\alpha}^{i}=E_{\alpha}\cap\Omega_{i}$ and $E_{\beta}^{i}=E_{\beta}\cap\Omega_{i}$. Now we consider the following two Dirichlet problems of harmonic functions:
$\Delta u=0,\quad u=0\,\,\mbox{on}\,\,\partial\Omega_{i}\setminus\partial E_{\alpha}^{i}\quad\mbox{and}\quad u=1\,\,\mbox{on}\,\,\partial E_{\alpha}^{i};$
@@
$\Delta u=0,\quad u=0\,\,\mbox{on}\,\,\partial\Omega_{i}\setminus\partial E_{\beta}^{i}\quad\mbox{and}\quad u=1\,\,\mbox{on}\,\,\partial E_{\beta}^{i}.$
Thus we can obtain two sequences of harmonic functions, denoted as $\{u_{\alpha}^{i}\}$ and $\{u_{\beta}^{i}\}$. Since $(M,g)$ enjoys a Sobolev inequality (1.8), by the same arguments as in [58] and [26] we can see that there exists two harmonic functions $u_{\alpha}$ and $u_{\beta}$ on $(M,g)$ such that by neglecting two subsequences $u_{\alpha}^{i}$ and $u_{\beta}^{i}$ converge in the sense of $C^{k}$ ($k\geq 2$) to $u_{\alpha}$ and $u_{\beta}$ on any compact subset of $(M,g)$, where
$u_{\alpha}(x)\to 1\quad\mbox{as}\,\,x\in E_{\alpha}\to\infty\quad\mbox{and}\quad u_{\alpha}(x)\to 0\quad\mbox{as}\,\,x\in M\setminus E_{\alpha}\to\infty;$
and
$u_{\beta}(x)\to 1\quad\mbox{as}\,\,x\in E_{\beta}\to\infty\quad\mbox{and}\quad u_{\beta}(x)\to 0\quad\mbox{as}\,\,x\in M\setminus E_{\beta}\to\infty.$
Obviously, $u_{\alpha}$ and $u_{\beta}$ are linearly independent in the linear space spanned by bounded harmonic functions on $(M,g)$. In other words, the dimension of $H^{\infty}(M)$ is at least two, i.e., $\dim(H^{\infty}(M))\geq 2$. For more details we refer to [26] and [58].
Proof of Theorem 1.6: We prove by contradiction. If $(M,g)$ has at least two ends, then, by the assumption the Sobolev constant of $(M,g)$ is positive, we can take the same argument as in [58] or [11] to conclude that there exists a nonconstant, bounded, and positive harmonic function on $(M,g)$. For more details we refer to Theorem B in [58] and Corollary 4.3 in [26]. However, Corollary 1.4 tells us that there exists a positive constant $C(n,\beta^{*})$ depending on $n$ and $\beta^{*}$ such that, if

### v2

@@ Abstract.
As topological applications of main results(see Corollory 1.8) we show that for a complete noncompact Riemannian manifold on which the Sobolev inequality (1.8) holds true, $\dim(M)=n\geq 3$ and ${\rm Ric}(x)\geq 0$ outside some geodesic ball $B(o,R_{0})$, there exists a positive constant $C(n)$ depending only on $n$ such that, if
@@ 1. Introduction
The renowned Cheng-Yau’s logarithm gradient estimate showed that when $p=2$, any solution bounded from above or below to (1.4) is a constant provided the Ricci curvature of the Riemannian manifold is nonnegative (cf. [16]). Kotschwar-Ni [35] proved that any positive $p$-harmonic function on a complete Riemannian manifold with nonnegative sectional curvature is a positive constant. Subsequently, this result was verified to remain correct by Wang and Zhang [59] under the condition of nonnegative Ricci curvature, where the authors applied the Nash-Moser iteration technique and Saloff-Coste’s Sobolev inequality (cf. [50, Theorem 3.1]) to study the gradient estimates of equation (1.4). This gradient estimate was later refined to be sharp by Sung and Wang [56].
@@
Especially, in the case $p=2$, $a>0$ and $q\in(-\infty,\,\frac{n+2}{n-2})$, Lu [41] established the Cheng-Yau type logarithm gradient estimate for positive solutions to Lane-Emden equation (1.2) on a complete Riemannian manifold with Ricci curvature bounded from below.
@@
It is well-known that (1.7) holds true for a large class of Cartan-Hadamard manifolds on which the Poincaré inequalities hold, for details we refer to [33]. We also notice that if the $\chi$-type Sobolev inequality holds on $(M,g)$ with $\mathrm{dim}(M)\geq 3$, then $\chi\leq n/(n-2)$ (see Theorem 2.1). Hence, we assume $\chi\in(1,n/(n-2)]$ when $\mathrm{dim}(M)\geq 3$ in the rest of this article. Clearly, (1.7) is just the well-known Sobolev inequality if $\chi={n}/(n-2)$, i.e.
@@
which holds true for a large class of complete Riemannian manifolds. Indeed, significant efforts have recently been devoted to studying optimal Sobolev inequalities on Riemannian manifolds (see the survey [27] and its references).
@@
It was first observed by the first named author [60, Proposition 2.4], and later by Carron (cf. [12]) and Akutagawa (cf. [1]) independently that if the Sobolev constant of a Riemannian manifold $(M^{n},g)$ is positive, then the volume of the geodesic ball of radius $R$ is larger than $CR^{n}$, where the constant $C$ depends only on $n$ and the Sobolev constant.
@@
    In a certain sence, these two theorems can be seen as an effective version of the classical Liouville results due to Gidas-Spruck [29] and Serrin-Zou [53]. To see this, let $g_{\epsilon}$ be a sequence of metrics on $\mathbb{R}^{n}$ with $g_{\epsilon}$ equals to the Euclidean metric outside some compact set of $\mathbb{R}^{n}$ and with $g_{\epsilon}$ converges to the Euclidean metric in $C^{2}$ sense as $\epsilon\rightarrow 0$. Then the Positive Mass Theorem tells us that $g_{\epsilon}$ can not be of nonnegative Ricci curvature. Since the Sobolev inequality only depends on the $C^{0}$ property of the metric. Our results indicate that the classical Liouville property also holds if the deformed metric $g_{\epsilon}$ satisfied
@@
Now, we turn to recalling the topological properties of a noncompact complete manifold related to the notion of end. “An end $E$ of a manifold $M$ is an unbounded component of the complement of some compact subset $D$ of $M$. In this case, we say that $E$ is an end corresponding to DD”.
For a noncompact complete manifold with nonnegative Ricci curvature outside a compact set, in 1991 Cai [9] showed that such a manifold is of finitely many ends. Later, in 1995 Cai, Colding and Yang [10] discussed the gap phenomenon for ends of such a class of manifold, concretely, they proved the following theorem:
@@
On the other hand, using harmonic function theory to study the number of ends has a long history (see for example, [11, 26, 36]). Especially, if the Sobolev constant of $(M,g)$ is positive and $\dim(M)\geq 3$, the first named author has ever used harmonic function theory to prove that such a manifold is of finitely many ends and the dimension of linear space spanned by bounded harmonic functions on $(M,g)$ equals the number of ends (cf. [60]).
By the volume estimate established in Theorem 1.1 we can employ the method adopted in [26] and [60] to show the following theorem.
Theorem 1.6.
Let $(M,g)$ be a complete noncompact Riemannian manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$. If $(M,g)$ has at least $k$ ends, then $\dim(H^{\infty}(M))\geq k$.
As a geometric application of Corollory 1.4 and Theorem 1.6, we obtain the following gap theorem for ends.
@@
The rest of our paper is organized as follows. In section 2, we will give the volume estimate of a geodesic ball $B_{R}$ in $(M,g)$ on which a $\chi$-type Sobolev inequality holds. Section 3 is devoted to a meticulous estimate of $\mathcal{L}\left(|\nabla\log v|^{2\alpha}\right)$ (the explicit definition of the operator $\mathcal{L}$is given in (3.3)). The proofs of our main results reveal that, by selecting an appropriate parameter $\alpha$, we can establish effective integral estimates for the gradient of positive solutions to equation (1.1). In particular, when $\rm Ric$$_{-}$ satisfies the condition stated in Theorem 1.2 or Theorem 1.5, we obtain a $L^{\theta\chi}$ bound of $|\nabla\log v|$. The crucial thing is this bound only depends on the volume and the radius of the geometric ball. With this integral bound in hand, we then apply the Nash-Moser iteration scheme to complete the proof of Theorem 1.5. In section 4, we continue to prove Theorem 1.3 by choosing suitable auxiliary functions. In Section 5 we provide the proof of Theorem 1.6 and Theorem 1.7.
@@ 2. Volume estimate
Hence, we conclude from the above inequality that
@@
$\lim_{m\rightarrow\infty}\alpha(m)=\frac{\chi}{\chi-1}\quad\quad\text{and}\quad\quad\lim_{m\rightarrow\infty}\beta(m)=\frac{\chi(2\chi-1)}{(\chi-1)^{2}}.$
@@ 5. Geometric applications
Now we want to make use of some conclusions obtain in the previous sections to study the topological properties of a manifold which enjoy a Sobolev inequality. As mentioned in the previous, many manifolds, for instances Riemannian manifold with ${\rm Ric}\geq 0$ and $0<\mathrm{AVR}_{g}\leq 1$, minimal hypersurfaces [43] and Ricci shrinker [39], support Sobolev inequalities. In particular, Brendle [7] obtained the sharp Sobolev inequality on complete noncompact Riemannian manifolds with nonnegative Ricci curvature using the so-called ABP method, addressing an open question posed by Cordero-Erausquin, Nazaret, and Villani [21] for such manifold. Balogh and Kristály [3] provided an alternative proof of Brendle’s rigidity result and confirmed that the Sobolev constant is sharp. Concretely, their results can be stated as follows:
and $\omega_{n}$ is the volume of the unit ball in $\mathbb{R}^{n}$ and $\mathcal{S}\left(\mathbb{R}^{n}\right)$ is the best Sobolev constant in $\mathbb{R}^{n}$.
Using harmonic function theory to study geometric and topological properties of manifolds has a long history, for more details we refer to [36, 37, 60] and references therein. Here, firstly we need to recall the definition of end in a complete noncompact manifold:
Definition.
Let $(M,g)$ be a complete noncompact Riemannian manifold. We say that $M$ has $k$ ends $E_{1},\cdots,E_{k}$, if there exists a bounded open set $U$ so that
$M\setminus\bar{U}=\cup_{i=1}^{k}E_{i}$
with each $E_{i}$ being a noncompact connected component of $M\setminus\bar{U}$. Usually, we also call $E_{1},\cdots,E_{k}$ as the ends corresponding to $U$.
Denote usually the linear space spanned by bounded harmonic functions on $(M,g)$ by $H^{\infty}(M)$. Now let’s recall some previous conclusions. The first named author of this paper has ever shown the following:
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold with $\dim(M)\geq 3$, Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)>0$ and ${\rm Ric}(M)\geq 0$ outside some compact subset. Then $M$ has only finitely many ends $E_{1},E_{2},\cdots,E_{k}$ and $\dim H^{\infty}(M)=k$.
@@
As an extension of Theorem 5.1, we can take almost the same arguments as in [26] and [60] to conclude the following theorem, i.e., Theorem 1.6.
Theorem 5.2.
Let $(M,g)$ be a complete noncompact Riemannian manifold with Sobolev constant $\mathbb{S}_{\chi}(M)>0$. If $(M,g)$ has at least $k$ ends, then $\dim(H^{\infty}(M))\geq k$.
Proof.
@@
Actually, by taking almost the same arguments as in [26] and [60], we can show the theorem. For the sake of completeness, here we give the routine to construct bounded positive harmonic functions on such a manifold $(M,g)$ which has at least $k$ ends and the Sobolev constant $\mathbb{S}_{\chi}(M)$ of $(M,g)$ is of a positive lower bound.
We let $\Omega_{i}$ ($i=1,2,\cdots$) be an exhaustion of $(M,g)$, i.e., each $\Omega_{i}$ is an open domain contained in $M$ and $\bar{\Omega}_{i}$ is compact, $\bar{\Omega}_{i}\subset\Omega_{i+1}$ for every $i\geq 1$, and $\cup_{i=1}^{\infty}\Omega_{i}=M$. For $i\geq i_{0}$ and $i_{0}$ sufficiently large, let
$M\setminus\Omega_{i}=\cup^{s}_{j=1}E^{(i)}_{j},\quad s\geq k$
be the disjoint union of connected components. Denote the $k$ ends of $(M,g)$ by $E_{1}^{(i_{0})},\cdots,E_{k-1}^{(i_{0})}$ and $E_{k}^{(i_{0})}$, and for the sake of convenience, by $E_{1},\cdots,E_{k}$ without confusions.
@@
Moreover, denote $E_{j}^{i}=E_{j}\cap\Omega_{i}$. Now we consider the following Dirichlet problems of harmonic functions:
$\Delta u_{j}^{i}=0,\quad u=0\,\,\mbox{on}\,\,\partial\Omega_{i}\setminus\partial E_{j}^{i}\quad\mbox{and}\quad u_{j}^{i}=1\,\,\mbox{on}\,\,\partial E_{j}^{i},$
where $j=1,\cdots,k$. Thus we can obtain $k$ sequences of harmonic functions, denoted as $\{u_{j}^{i}\}$ ($j=1,2,\cdots,k$).
As done in [26], we let $\phi_{j}(x)$ be a real smooth function on $(M,g)$ such that
$\phi_{j}(x)\equiv 1\quad\mbox{on}\,\,E_{j}^{(i_{0})}\quad\mbox{and}\quad\phi_{j}(x)\equiv 0\quad\mbox{in}\,\,M\setminus(\Omega_{i_{0}}\cup E_{j}^{(i_{0})});$
and then define $\rho_{j,i}$ on $\Omega_{i}$ by
$\rho_{j,i}=|u^{i}_{j}-\phi_{j}|\quad\mbox{for}\,\,i\geq i_{0}.$
It is easy to see that $|\Delta\phi_{j}|\in L^{p}$ for any $p>0$ and in the sense of distribution (see [26])
$\Delta\rho_{j,i}\geq-|\Delta\phi_{j}|.$
We choose a test function $\rho_{j,i}^{\alpha-1}$ ($\alpha>1$) to multiply the above equation and integrate on $\Omega_{i}$
$\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}\Delta\rho_{j,i}\geq-\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
Then we have
$\frac{4(\alpha-1)}{\alpha^{2}}\int_{\Omega_{i}}\left|\nabla\rho_{j,i}^{\frac{\alpha}{2}}\right|^{2}\leq\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
Since $(M,g)$ enjoys a Sobolev inequality as follows
$\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv,$
we obtain
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq\frac{\alpha^{2}}{4(\alpha-1)}\mathbb{S}_{\chi}(M)^{-1}\int_{\Omega_{i}}\rho_{j,i}^{\alpha-1}|\Delta\phi_{j}|.$
Applying the Hölder inequality we can see
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\chi}}\leq C\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{\alpha-1}{\alpha\chi}}\left(\int_{\Omega_{i}}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$
It follows that
$\left(\int_{\Omega_{i}}\rho_{j,i}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}\leq C\left(\int_{\Omega_{i}}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}\right)^{\frac{\alpha\chi-\alpha+1}{\alpha\chi}}.$ (5.1)
By the same arguments as in [60] and Corollary 4.3 of [26] we can see that there exists $k$ harmonic functions $u_{1},\cdots,u_{k-1}$ and $u_{k}$ on $(M,g)$ such that by neglecting some subsequences $u_{1}^{i},\cdots,u_{k-1}^{i}$ and $u_{k}^{i}$ converge respectively in the sense of $C^{k}$ ($k\geq 2$) to $u_{1},\cdots,u_{k-1}$ and $u_{k}$ on any compact subset of $(M,g)$, where $j=1,\cdots,k$, i.e.,
$\lim_{i\to\infty}\|u^{i}_{j}-u_{j}\|_{C^{2}(\mathcal{C})}=0\quad\quad j=1,\cdots,k$
where $\mathcal{C}\in M$ is any compact subset. In fact, we can also use directly Cheng-Yau’s local gradient estimate [16] to show the convergence of these sequences $\{u^{i}_{j}\}$ ($j=1,\cdots,k;\,\,i=1,2,\cdots$).
Now, we claim that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent in the linear space spanned by bounded harmonic functions on $(M,g)$, i.e., $\dim(H^{\infty}(M))\geq k$.
In order to prove the above assertion, firstly we note that by maximum principle there hold true $0<u_{j}(x)<1$ for any $x\in M$, where $j=1,2,\cdots,k$. Then, we let $i\to\infty$ in (5.1)and make use of a diagonal technique to obtain
$\left(\int_{E_{j}}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\quad\mbox{and}\quad\quad\left(\int_{E_{l}}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}<\infty\quad\mbox{for}\,\,l\neq j,$ (5.2)
since we have
$\int_{M}|\Delta\phi_{j}|^{\frac{\alpha\chi}{\alpha\chi-\alpha+1}}<\infty.$
Therefore, it is easy to verify from the above integral estimates that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are not constant functions, since ${\rm{vol}}(E_{j})=\infty$ for each $j=1,\cdots,k$.
The remaining is to show that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent. Since for any geodesic ball $B_{r}\subset M$ we have the volume estimate
$\mathrm{Vol}\left(B_{r}\right)\geq C\left(\chi,\mathbb{S}_{\chi}(M)\right)r^{\frac{2\chi}{\chi-1}},$
this tells us that $\mathrm{Vol}(B_{r}(x))$ is of a uniform lower bound. Hence, noting $\chi>1$ and $\alpha>1$ we have
$\frac{1}{\mathrm{Vol}(B_{r}(x))}\int_{B_{r}(x)}u_{j}\leq\left(\frac{1}{\mathrm{Vol}(B_{r}(x))}\int_{B_{r}(x)}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}.$
It follows from (5.2) that for $B_{r}(x)\subset E_{i}$, where $r>0$ is a fixed number, and $j\neq i$ there hold true
$\lim_{x\in E_{i}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{j}=\lim_{x\in E_{i}\to\infty}\mathrm{Vol}(B_{r}(x))^{-\frac{1}{\alpha\chi}}\left(\int_{B_{r}(x)}u_{j}^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}=0$
@@
$\lim_{x\in E_{j}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{j})=\lim_{x\in E_{j}\to\infty}\mathrm{Vol}(B_{r}(x))^{-\frac{1}{\alpha\chi}}\left(\int_{B_{r}(x)}(1-u_{j})^{\alpha\chi}\right)^{\frac{1}{\alpha\chi}}=0.$
If there holds
$\sum_{j=1}^{k}a_{j}u_{j}(x)\equiv 0\quad\mbox{on}\,\,M,$
where $a_{j}$ ($j=1,2,\cdots,k$) are some real numbers, then we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}\sum_{j=1}^{k}a_{j}u_{j}=0,$
which implies
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}a_{1}u_{1}=0.$
We conclude that $a_{1}=0$. Otherwise, we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=0.$
On the other hand, by the integral estimates (5.2) we have
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})=0.$
Hence, it follows
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})+\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=0.$
However, we have that for any $B_{r}(x)\subset E_{1}$ there holds
$1=\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}1=\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}\{(1-u_{1})+u_{1}\}$
which implies
$\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}(1-u_{1})+\lim_{x\in E_{1}\to\infty}\mathrm{Vol}(B_{r}(x))^{-1}\int_{B_{r}(x)}u_{1}=1.$
This is a contradiction. Similarly, we can also show that $a_{2}=0,\cdots,a_{k-1}=0$ and $a_{k}=0$. This indicates that $u_{1},\cdots,u_{k-1}$ and $u_{k}$ are linearly independent, hence we have shown
$\dim(H^{\infty}(M))\geq k.$
Thus, we finish the proof. ∎
Remark 2.
In fact, Theorem 5.2 can be extended to the following case $(M,g)$ enjoys a more general Sobolev inequality:
$\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}(|\nabla f|^{2}+Rf^{2})dv,\quad\mbox{for any}\,\,f\in C_{0}^{\infty}(M),$
which appears in the geometry of submanifolds and Ricci solitons (see [39, 43]). Here $R(x)\geq 0$ is usually a smooth function related to the curvature. Obviously, it is easy to verify the inequality reduces to
$\mathbb{C}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv,\quad\mbox{for any}\,\,f\in C_{0}^{\infty}(M),$
provided $\|R\|_{L^{\frac{\chi}{\chi-1}}}<\mathbb{S}_{\chi}(M)$. Particularly, such inequality with $\chi=\frac{n}{n-2}$ holds true in a complete minimal hypersurface in an Euclidean space (see [7, 43]).
Proof of Theorem 1.7: We prove this theorem by contradiction. If $(M,g)$ has at least two ends, then, by the assumption the $\frac{n}{n-2}$-type Sobolev constant $\mathbb{S}_{\frac{n}{n-2}}(M)$ is positive and Theorem 1.6 we conclude that there exists a nonconstant, bounded, and positive harmonic function on $(M,g)$ (see also Theorem B in [60] and Corollary 4.3 in [26] for more details). However, Corollary 1.4 tells us that there exists a positive constant $C(n,\beta^{*})$ depending on $n$ and $\beta^{*}$ such that, if